2x3 – 5xy - y2 = 3 dy Find da Choose 1 answer: 6x2 - 57 5x+2y 6x - 2y 5 3 + 2y 6x2-5 6x2-57 2y - 5
Answer #1

223 -- bzy- y2- 3 d da (223 - 57y - y2) = d dz (3) (Xy) – 2y dy =0 6x² - 5 år de Cay) > 622 - 5[zas ty dx zyky da =0 => 6x² -